# Jailbreaking and Mitigation of Vulnerabilities in Large Language Models

### v1

@@ I Introduction
Researchers have indicated that LLMs can be manipulated to provide instructions for illegal activities such as drug synthesis, bomb-making, and money laundering [8]. Other studies have demonstrated the effectiveness of persuasive language, based on social science research, in jailbreaking LLMs to generate harmful content [9]. Multilingual prompts can exacerbate the impact of malicious instructions by exploiting linguistic gaps in safety training data, leading to high rates of unsafe output [10]. These attacks show the fragility of current safety alignment techniques and the need for more robust defenses [11]. Qi et al. [12] demonstrated that fine-tuning aligned LLMs, even with benign data, can compromise safety.
@@ II-B Prompt Engineering
Prompt engineering is the process of designing the input text, or prompt, given to an LLM to elicit the desired output [6][35]. It plays a crucial role in enhancing LLM performance and ensuring safety by providing context, specifying the task, and guiding the model’s behavior. Effective prompts significantly improves the accuracy, relevance, and creativity of the generated text, while also mitigating the risk of harmful or biased outputs. Prompt engineering involves a variety of techniques, ranging from simple instructions to more complex strategies that fully utilize the LLM’s capabilities. Zero-shot prompting involves providing a task description without any examples [36], while few-shot prompting includes a few examples to guide the model [36]. Chain-of-thought prompting encourages the LLM to generate a step-by-step reasoning process before providing the final answer [35], while tree-of-thought prompting expands on this by exploring multiple reasoning paths [35]. Role prompting assigns a specific role or persona to the LLM [35], whereas instruction prompting provides explicit instructions to generate the desired output format or content. Bespoke prompt engineering enhances LLM safety and mitigates risks, which involves designing prompts that instruct the LLM to avoid generating harmful or unethical content explicitly, respect diverse perspectives, and adhere to established ethical guidelines. For example, prompts may instruct the LLM to avoid hate speech, consider cultural sensitivities, or prioritize factual accuracy over creative storytelling. In some cases, prompts can remind the LLM of its safety guidelines and responsibilities, serving as a form of self-regulation [37].
@@ II-C Jailbreaking
The rise of online communities sharing jailbreak prompts has hugely escalated the threat levels. These communities collaborate to discover vulnerabilities, refine attacks, and bypass new defenses [34][11]. The rapid evolution and growing sophistication of jailbreaking highlight the need for continuous development of robust defenses. The shift to dedicated prompt-aggregation websites signals a trend towards more organized and sophisticated jailbreaking [7].
@@ III-C1 Visual Jailbreaking
ImgTrojan: ImgTrojan poisons the training data by replacing original image captions with malicious jailbreak prompts [39]. When the poisoned image is presented to the model, the embedded prompt triggers the generation of harmful content. This attack underscores the risk of backdoor vulnerabilities in multimodal LLMs.
@@
FigStep: FigStep converts harmful text into images using typography, bypassing the safety mechanisms in the MLLM’s text module [21]. It gaps in safety alignment between visual and textual modalities, therefore achieving high success rates against various open-source VLMs.
@@ IV-C1 Collaborative Filtering
Collaborative filtering involves using multiple LLM agents with different roles and perspectives to analyze and filter out harmful responses. This approach benefits from the combined knowledge and reasoning abilities of multiple LLMs, making it more difficult for attackers to bypass the defenses. An example is the AutoDefense framework, which assigns different roles to LLM agents and uses them to collaboratively analyze and filter harmful outputs, enhancing the system’s robustness against jailbreaking attacks while maintaining normal performance for benign queries [69]. However, this approach also requires careful coordination and communication between the agents to ensure effective collaboration and avoid potential conflicts or inconsistencies in their decisions.
@@ IV-D3 Safety-Aware Decoding
Safety-aware decoding modifies the LLM decoding process to prioritize safe outputs and mitigate jailbreak attacks. SafeDecoding amplifies the probabilities of safety disclaimers in generated text while reducing the probabilities of token sequences linked to jailbreak objectives [74]. This approach uses safety disclaimers present in potentially harmful outputs, enabling the decoder to prioritize them and reduce harmful content. However, this method may make the model overly cautious, causing it to refuse responses to benign prompts that contain sensitive keywords.
@@ VI Research Gaps and Future Directions
To illustrate these vulnerabilities, we conducted some testing with the latest GPT-4o model and Perplexity Pro, as shown in Figure 3. Our experiments revealed that, despite integrated safety measures, both models were susceptible to simple yet carefully crafted prompts, resulting in the unintended disclosure of system instructions. Notably, a simple translation request originating from an educational scenario in Traditional Chinese, bypassed its safety mechanisms and prompted the release of privileged model-level information.
@@ VI-G2 Complex Interplay Between Capabilities and Safety
Further research is required to understand the relationship between LLM capabilities and safety. [58] identified two failure modes of safety training—competing objectives and mismatched generalization—highlighting the need for advanced safety mechanisms that match LLM sophistication.
@@ VII-B Implications for Research and Practice
The findings underscore an urgent need to rethink how LLMs are developed and deployed. Simply scaling models or applying superficial safety measures is insufficient. [10] found that multilingual prompts can exacerbate malicious instructions, emphasizing the need for safeguards that cover diverse linguistic contexts.
@@ VII-B3 Utilizing LLM Capabilities for Defense
The capabilities that make LLMs vulnerable can also be used for defense. [84] proposed SELFDEFEND, using the LLM to detect harmful prompts and respond accordingly. [37] explored a self-reminder technique, reducing jailbreak success rates by encapsulating queries in responsible system prompts. Further research should make use of LLMs’ strengths in language understanding to create adaptive defense mechanisms.
@@ VII-C Path Forward
The findings of this review emphasize the need for a collaborative effort to address LLM security and safety challenges. As LLMs become more powerful and integrated into critical applications, the risks of misuse increase. We encourage the AI community to prioritize research on robust alignment, effective defense mechanisms, and comprehensive evaluation frameworks for responsible deployment. Collaboration between researchers, industry, policymakers, and the public is crucial for establishing ethical guidelines and best practices in using these powerful technologies. By working together, we can mitigate risks and ensure the beneficial impact of LLMs on society.

### v2

@@ I Introduction
Researchers have indicated that LLMs can be manipulated to provide instructions for illegal activities such as drug synthesis, bomb-making, and money laundering [8]. Other studies have demonstrated the effectiveness of persuasive language, based on social science research, in jailbreaking LLMs to generate harmful content [9]. Multilingual prompts can exacerbate the impact of malicious instructions by exploiting linguistic gaps in safety training data, leading to high rates of unsafe output [10]. These attacks highlight the limitations of current safety alignment techniques and underscore the need for more robust defenses [11]. Qi et al. [12] demonstrated that fine-tuning aligned LLMs, even with benign data, can compromise safety.
@@ II-B Prompt Engineering
Prompt engineering is the process of designing the input text, or prompt, given to an LLM to elicit the desired output [6][35]. It plays a crucial role in enhancing LLM performance and ensuring safety by providing context, specifying the task, and guiding the model’s behavior. Effective prompts improve the accuracy, relevance, and creativity of the generated text, while reducing the risk of harmful or biased outputs. Prompt engineering involves a variety of techniques, ranging from simple instructions to more complex strategies that fully utilize the LLM’s capabilities. Zero-shot prompting involves providing a task description without any examples [36], while few-shot prompting includes a few examples to guide the model [36]. Chain-of-thought prompting encourages the LLM to generate a step-by-step reasoning process before providing the final answer [35], while tree-of-thought prompting expands on this by exploring multiple reasoning paths [35]. Role prompting assigns a specific role or persona to the LLM [35], whereas instruction prompting provides explicit instructions to generate the desired output format or content. Bespoke prompt engineering enhances LLM safety and mitigates risks, which involves designing prompts that instruct the LLM to avoid generating harmful or unethical content explicitly, respect diverse perspectives, and adhere to established ethical guidelines. For example, prompts may instruct the LLM to avoid hate speech, consider cultural sensitivities, or prioritize factual accuracy over creative storytelling. In some cases, prompts can remind the LLM of its safety guidelines and responsibilities, serving as a form of self-regulation [37].
@@ II-C Jailbreaking
The rise of online communities sharing jailbreak prompts has significantly increased threat levels. These communities collaborate to discover vulnerabilities, refine attacks, and bypass new defenses [34][11]. The rapid evolution and growing sophistication of jailbreaking highlight the need for continuous development of robust defenses. The shift to dedicated prompt-aggregation websites signals a trend towards more organized and sophisticated jailbreaking [7].
@@ III-C1 Visual Jailbreaking
ImgTrojan: ImgTrojan poisons the training data by replacing original image captions with malicious jailbreak prompts [39]. When the poisoned image is presented to the model, the embedded prompt triggers the generation of harmful content. This attack highlights the severity of backdoor vulnerabilities in multimodal LLMs.
@@
FigStep: FigStep converts harmful text into images using typography, bypassing the safety mechanisms in the MLLM’s text module [21]. It exploits gaps in safety alignment between visual and textual modalities, achieving high success rates against various open-source VLMs.
@@ IV-C1 Collaborative Filtering
Collaborative filtering involves using multiple LLM agents with different roles and perspectives to analyze and filter out harmful responses. This approach leverages the combined knowledge and reasoning abilities of multiple LLMs, thereby increasing the difficulty for attackers to bypass defenses. An example is the AutoDefense framework, which assigns different roles to LLM agents and uses them to collaboratively analyze and filter harmful outputs, enhancing the system’s robustness against jailbreaking attacks while maintaining normal performance for benign queries [69]. However, this approach also requires careful coordination and communication between the agents to ensure effective collaboration and avoid potential conflicts or inconsistencies in their decisions.
@@ IV-D3 Safety-Aware Decoding
Safety-aware decoding modifies the LLM decoding process to prioritize safe outputs and mitigate jailbreak attacks. SafeDecoding amplifies the probabilities of safety disclaimers in generated text while reducing the probabilities of token sequences linked to jailbreak objectives [74]. This approach uses safety disclaimers present in potentially harmful outputs, enabling the decoder to prioritize them and reduce harmful content. However, this method may lead the model to become excessively cautious, resulting in refusals of benign prompts containing sensitive keywords.
@@ VI Research Gaps and Future Directions
To illustrate these vulnerabilities, we conducted preliminary tests using the latest GPT-4o model and Perplexity Pro, as shown in Figure 3. Our experiments revealed that, despite integrated safety measures, both models were susceptible to simple yet carefully crafted prompts, resulting in the unintended disclosure of system instructions. Notably, a simple translation request originating from an educational scenario in Traditional Chinese, bypassed its safety mechanisms and prompted the release of privileged model-level information.
@@ VI-G2 Complex Interplay Between Capabilities and Safety
Further research is necessary to better understand the relationship between LLM capabilities and safety. [58] identified two failure modes of safety training—competing objectives and mismatched generalization—highlighting the need for advanced safety mechanisms that match LLM sophistication.
@@ VII-B Implications for Research and Practice
The findings underscore an urgent need to rethink how LLMs are developed and deployed. Merely scaling models or applying surface-level safety measures remains insufficient. [10] found that multilingual prompts can exacerbate malicious instructions, emphasizing the need for safeguards that cover diverse linguistic contexts.
@@ VII-B3 Utilizing LLM Capabilities for Defense
The capabilities that make LLMs vulnerable can also be used for defense. [84] proposed SELFDEFEND, using the LLM to detect harmful prompts and respond accordingly. [37] explored a self-reminder technique, reducing jailbreak success rates by encapsulating queries in responsible system prompts. Further research should leverage LLMs’ strengths in language understanding to develop adaptive defense mechanisms.
@@ VII-C Path Forward
The findings of this review underscore the importance of collaborative efforts to address LLM security and safety challenges. As LLMs become more powerful and integrated into critical applications, the risks of misuse increase. We encourage the AI community to prioritize research on robust alignment, effective defense mechanisms, and comprehensive evaluation frameworks for responsible deployment. Collaboration between researchers, industry, policymakers, and the public is crucial for establishing ethical guidelines and best practices in using these powerful technologies. By working together, we can mitigate risks and ensure the beneficial impact of LLMs on society.